# Ion Coulomb crystals: an exotic form of condensed matter

### v1

@@ IV Concluding remarks and outlook
In contrast with electronic systems in condensed matter, ion Coulomb crystals give access to a regime that is dominated by interactions, rather than the kinetic energy of the particles. Their equilibrium and out-of-equilibrium dynamics is strongly shaped by the long-range potential, where the dominant energy contribution is non-additive, thus breaking crucial principals of textbook statistical mechanics, such as ensemble equivalence [45]. This makes ion-Coulomb crystals a prominent laboratory for testing theoretical hypotheses formulated for astrophysical and plasma systems [45]. Moreover, the experimental progress opens the possibility for accessing regimes so far unexplored that can shed light on the onset of equilibrium [73] and on quantum phases [72] in the presence of long-range, non-additive forces.
@@
Exotic quantum phases and critical phenomena are, for instance, the zigzag and buckling structural phase transitions, as well as the topological features of frustrated geometries, such as the Aubry and commensurate-incommensurate transition in ion chains. These are also examples of how discrete internal degrees of freedom, such as spins, are naturally encoded in the structure of a low dimensional crystal. This idea could be generalized to multilayer crystals where the layer index is a discrete degree of freedom that serves as an effective spin, mapping the dynamics to models analogous to antiferromagnetism in frustrated geometries. The long-range nature of topological defects, such as dislocations and kinks, is only partially characterized in one dimension, and unexplored in higher dimensions. One question is how the coupling between lattice vibrations and topological defects (dislocations, disclinations, kinks, vortices, etc) might give rise to metastable structures and novel phases in frustrated geometries with long-range Coulomb forces. This can be seen as an exotic realization of quantum magnetism in dynamic Wigner lattices, see [233] for a related proposal. It further holds the promise for simulating dynamical features of lattice-field theory with trapped ions [15, 208, 17]. It would be interesting, for example, to investigate how the influence of the coupling to the lattice vibrations and defects affects the nature of the magnetic phases (and thus of the layer ordering) as well as the transitions between them [289, 274, 275, 2].
@@
One peculiar aspect of these systems is that exchange interaction and quantum statistics are very small and generally not measurable. Yet, an ingeneous experiment measured the Aharonov-Bohm effect in the macroscopic tunneling of a chain of three ions realizing a quantum rotor [235]. Protocols have been proposed for revealing quantum statistics in ion Coulomb crystals [264], indicating an avenue for unveiling these phenomena with larger crystals.
The progress reported in this review focused on fundamental aspects of ion Coulomb crystals, while we purposely did not discuss the formidable progress in using these systems for designing target quantum dynamics for quantum metrology, quantum simulation, quantum sensing, and quantum computing. Yet, the advances we reported have been going hand-in-hand with the development of these applications, giving rise to a cross fertilization of methods and insights. This progress allows us to anticipate that the scientific and technical applications of ion Coulomb crystals will continue to grow. Several of these applications are rooted in the fact that the internal states of the individual ions constituting the Coulomb crystals are only very weakly perturbed by the presence of the other ions. This makes ion Coulomb crystals highly interesting platform for high-resolution spectroscopy, sensing, and metrology [207]. Example applications in quantum sensing include the development of extremely accurate optical atomic ion clocks [152, 35, 317], ultra-high resolution spectroscopy of molecular ions [57, 177], and weak electromagnetic field sensing [178, 271, 204, 124]. Most recently, a multi-ion clock composed of 115In+ and 172Yb+ ions reached the systematic uncertainty of $2.5\times 10^{-18}$ [137]. These quantum sensing capabilities can in turn be used as fundamental physics probes for tests of Lorentz invariance violation [206, 82], searches for variation in natural constants over time [267, 101] and for particles beyond the Standard Model [119, 299, 66, 81], and signatures of potential extensions to quantum mechanics [190].
Last but not least, the high degree of experimental control over trapped ion crystals has enabled ion Coulomb crystals to be one of the leading platform for quantum information processing, including quantum simulation [24, 37, 218] and quantum computing [59, 42, 229].
@@
We thus expect much more to come in this field, where atomic, molecular, optical, statistical, and, most recently, condensed-matter physics merge in verifying hypotheses and providing novel insights into the physics of Wigner crystallization with trapped ions. This progress opens new paths towards understanding the onset and stability of long-range interacting quantum structure and offers exciting perspectives for technological advances in the quantum control of complex systems.
@@ Acknowledgements.
The authors are indebted to our close collaborators Daniel Dubin, Shmuel Fishman, Herbert Walther, and David Wineland, with whom we had the privilege to work over the years on this fascinating subject. We further had the privilege to share many useful discussions and insights with a number of colleagues and collaborators. For reason of space we refrain from mentioning them all, and explicitly list our collaborators on ion Coulomb crystal work. (in alphabetical order) Matt Affolter, Francois Anderegg, Jens Baltrusch, Rainer Blatt, Joseph Britton, Andreas Buchleitner, Tommaso Calarco, Adolfo del Campo, Florian Cartarius, Allison Carter, Caroline Champenois, Oksana Chelpanova, Maria-Luisa Chiofalo, Cecilia Cormick, Gabriele De Chiara, J. Ignacio Cirac, Aurelien Dantan, C. Fred Driscoll, Jürgen Eschner, Thomas Fogarty, James Freericks, Jeffrey S. Hangst, Jan-Petter Hansen, Peter Horak, Liv Hornekær, X.-P. Huang, Wayne Itano, Endre Kajari, Martina Knoop, Jan Kiethe, Niels Kjærgaard, Haggai Landa, Thomas Lauprêtre, Ian D. Leroux, Jamir Marino, Thierry Matthey, Tanja E. Mehlstäubler, Raphael Menu, Travis Mitchell, Simone Montangero, Thomas O’Neil, Ramil Nigmatullin, Scott Parker, Martin Plenio, Dirk Reiß, Alex Retzker, Ana Maria Rey, Benni Reznik, Brian Sawyer, Tobias Schätz, Ferdinand Schmidt-Kaler, John P. Schiffer, Wolfgang Schleich, Athreya Shankar, Pietro Silvi, Cyrille Solaro, Bruno Taketani, Joseph Tan, Lars Timm, Vladan Vuletić, Jorge Yago-Malo, Christof Wunderlich. The authors are grateful to Daniel Dubin and to Jennifer Lilieholm for the critical reading of the manuscript.

### v2

@@ IV Concluding remarks and outlook
In contrast with electronic systems in condensed matter, ion Coulomb crystals give access to a regime that is dominated by interactions, rather than the kinetic energy of the particles. Their equilibrium and out-of-equilibrium dynamics is strongly shaped by the long-range potential, where the dominant energy contribution is non-additive, thus breaking crucial principles of textbook statistical mechanics, such as ensemble equivalence [49]. This makes ion Coulomb crystals a prominent laboratory for testing theoretical hypotheses formulated for astrophysical and plasma systems [49]. Moreover, the experimental progress opens the possibility for accessing regimes so far unexplored that can shed light on the onset of equilibrium [78] and on quantum phases [77] in the presence of long-range, non-additive forces.
@@
Exotic quantum phases and critical phenomena are, for instance, the zigzag and buckling structural phase transitions, as well as the topological features of frustrated geometries, such as the Aubry and commensurate-incommensurate transition in ion chains. These are also examples of how discrete internal degrees of freedom, such as spins, are naturally encoded in the structure of a low dimensional crystal. This paradigm could be generalized to multilayer crystals where the layer index is a discrete degree of freedom that serves as an effective spin, mapping the dynamics to models analogous to antiferromagnetism in frustrated geometries. The long-range nature of topological defects, such as dislocations and kinks, is only partially characterized in one dimension, and unexplored in higher dimensions. One question is how the coupling between lattice vibrations and topological defects (dislocations, disclinations, kinks, vortices, etc.) might give rise to metastable structures and novel phases in frustrated geometries with long-range Coulomb forces. This can be seen as an exotic realization of quantum magnetism in dynamic Wigner lattices, see [259] for a related proposal. It holds the promise for simulating dynamical features of lattice-field theory with trapped ions [16, 19, 232, 179]. It would be interesting, for example, to investigate how the influence of the coupling to the lattice vibrations and defects affects the nature of the magnetic phases (and thus of the layer ordering) as well as the transitions between them [324, 307, 2, 308].
@@
One peculiar aspect of these systems is that exchange interaction and quantum statistics are very small and generally not measurable. Yet, an ingenious experiment measured the Aharonov-Bohm effect in the macroscopic tunneling of a chain of three ions realizing a quantum rotor [262]. Protocols have been proposed for revealing quantum statistics in ion Coulomb crystals [296], indicating an avenue for unveiling these phenomena with larger crystals.
The progress reported in this review focused on fundamental aspects of ion Coulomb crystals, while intentionally leaving aside the remarkable advances in their use for engineered quantum dynamics in quantum metrology, quantum simulation [27, 39, 243], quantum sensing, and quantum computing [64, 44, 255]. Closely related developments include fundamental insights into tailoring correlations between ions with photons by means of optical elements and cavities [114, 53] and advanced detection schemes [241, 361, 295], which underpin powerful applications in quantum communication [96, 306]. Importantly, the advances reported here have evolved hand-in-hand with the development of these applications, leading to a strong cross-fertilization of methods and physical insights.
This interplay suggests that both the scientific reach and technological impact of ion Coulomb crystals will continue to grow. Many of these applications are rooted in the fact that the internal states of individual ions in a Coulomb crystal are only weakly perturbed by the presence of other ions, making these systems a highly interesting platform for high-resolution spectroscopy, sensing, and metrology [231]. Representative examples in quantum sensing include extremely accurate optical atomic ion clocks [164, 352, 222], ultra-high-resolution spectroscopy of molecular ions [62, 192], and precision sensing of weak electromagnetic fields [193, 303, 134, 228]. Notably, a recent multi-ion clock based on a linear chain of 4 115In+ ions sympathetically cooled by 8 172Yb+ ions demonstrated a fractional frequency instability of $9.2\times 10^{-16}$ at 1 second averaging time [149].
These quantum sensing capabilities can in turn be used as powerful probes of fundamental physics, enabling tests of Lorentz invariance [230, 88], searches for temporal variations of fundamental constants [299, 108], for particles beyond the Standard Model [129, 71, 334, 87], and for signatures of potential extensions to quantum mechanics [207].
@@
We thus expect much more to come in this field, where atomic, molecular, optical, statistical, and, most recently, condensed-matter physics merge in verifying hypotheses and providing novel insights into the physics of Wigner crystallization with trapped ions. This progress opens new paths towards understanding the onset and stability of long-range interacting quantum structures and offers exciting perspectives for technological advances in the quantum control of complex systems.
@@ Acknowledgements.
The authors are indebted to our close collaborators Daniel Dubin, Shmuel Fishman, Herbert Walther, and David Wineland, with whom we had the privilege to work over the years on this fascinating subject. We further had the privilege to share many useful discussions and insights with a number of colleagues and collaborators. For reasons of space we refrain from mentioning them all, and explicitly list our collaborators on ion Coulomb crystal work in alphabetical order: Matt Affolter, Francois Anderegg, Jens Baltrusch, Rainer Blatt, Joseph Britton, Andreas Buchleitner, Tommaso Calarco, Adolfo del Campo, Florian Cartarius, Allison Carter, Caroline Champenois, Oksana Chelpanova, Maria-Luisa Chiofalo, Cecilia Cormick, Gabriele De Chiara, J. Ignacio Cirac, Aurelien Dantan, C. Fred Driscoll, Jürgen Eschner, Thomas Fogarty, James Freericks, Jeffrey S. Hangst, Jan-Petter Hansen, Peter Horak, Liv Hornekær, X.-P. Huang, Wayne Itano, Endre Kajari, Martina Knoop, Jan Kiethe, Niels Kjærgaard, Haggai Landa, Thomas Lauprêtre, Dietrich Leibfried, Ian D. Leroux, Jamir Marino, Thierry Matthey, Tanja E. Mehlstäubler, Raphael Menu, Travis Mitchell, Simone Montangero, Thomas O’Neil, Ramil Nigmatullin, Scott Parker, Martin Plenio, Dirk Reiß, Alex Retzker, Ana Maria Rey, Benni Reznik, Brian Sawyer, Tobias Schätz, Ferdinand Schmidt-Kaler, John P. Schiffer, Wolfgang Schleich, Athreya Shankar, Pietro Silvi, Cyrille Solaro, Bruno Taketani, Joseph Tan, Lars Timm, Vladan Vuletić, Jorge Yago-Malo, Christof Wunderlich. The authors are grateful to Daniel Dubin and to Jennifer Lilieholm for the critical reading of the manuscript.